Q1.(A) Find equilibrium Price and quantity from the following demand and supply functions

        I.           D=40-4P                                                  S= 10+2P

(B)The demand function is Q= 20-5P estimate the elasticity of demand at P=2 and P=3

Equilibrium in market 



1
Expert's answer
2021-04-05T07:28:01-0400

Solution:

A.). Equilibrium price and quantity:

At equilibrium: D = S

40 – 4P = 10 + 2P

40 – 10 = 2P + 4P

30 = 6P

P = "\frac{30}{6} = 5"

Substitute for Q:

Q = 40 – 4P = 40 – 4(5) = 40 – 20 = 20

Equilibrium price = 5

Equilibrium quantity = 20

B.). Point price elasticity of demand:

At equilibrium: D = S

40 – 4P = 10 + 2P

40 – 10 = 2P + 4P

30 = 6P

P = 30/6 = 5

Substitute for Q:

Q = 40 – 4P = 40 – 4(5) = 40 – 20 = 20

Equilibrium price = 5

Equilibrium quantity = 20

B.). Elasticity of demand:

Q = 20 – 5P

When P = 2

Qd = 20 – 5 (2) = 20 – 10 = 10

When P = 3

Qd = 20 – 5 (3) = 20 – 15 = 5

Point price elasticity of demand (PED) = "\frac{ P}{ Q} \times\frac{\% \triangle Q}{ \%\triangle P} = \frac{ P}{ Q}\times\frac{ \triangle Q}{ \triangle P}"

Point price elasticity when P = 2,

First derive: Elasticity of demand

Elasticity of demand (Ed) = "\frac{ \triangle Q}{ \triangle P}" from the demand function:


Elasticity of demand (Ed) = "\frac{-5}{1} = - 5"


At P = 2, Q = 10


Point Price Elasticity of demand (PED) = "\frac{2}{10}\times - 5 = - 1"

PED = "-1"


At P = 3, Q = 5


Point Price Elasticity of demand (PED) = "\frac{3}{5}\times - 5 = - 3"

PED = "-3"



Point price elasticity for P = 2: = -1 and P = 3: = -3